The mean dying rate of mice administered to a given dosage of an experimental drug is 45with a variance of 25. What is the probability that among the mice, which received the drug, more than 50 will die? Assume normality



1
Expert's answer
2021-12-22T17:54:15-0500

Let "X=" the number of mice that died: "X\sim N(\mu, \sigma^2)."

Given "\mu=45, \sigma^2=25."


"P(X>50)=1-P(X\leq 50)"

"=1-P(Z\leq \dfrac{50-45}{\sqrt{25}})=1-P(Z\leq 1)"

"\approx0.158655"